# Shenzhen, China

### v4

@@ 0.G.1 Accuracy for time windows
EEGITNet 83.23 $\pm$ 2.77 65.67 $\pm$ 3.84 43.87 $\pm$ 3.78 7.79 $\pm$ 1.27 22.12 $\pm$ 2.88
EEGConformer 89.18 $\pm$ 2.88 75.49 $\pm$ 5.30 55.32 $\pm$ 6.88 11.29 $\pm$ 3.55 33.49 $\pm$ 6.78
ShallowFBCSPNet 84.29 $\pm$ 2.56 67.13 $\pm$ 3.61 46.07 $\pm$ 4.55 8.26 $\pm$ 2.95 25.87 $\pm$ 5.03
EEGNetV4 89.03 $\pm$ 2.86 75.40 $\pm$ 5.18 56.77 $\pm$ 6.60 13.29 $\pm$ 3.99 35.50 $\pm$ 7.14
B.D. 91.14 $\pm$ 2.45 79.59 $\pm$ 4.59 62.62 $\pm$ 6.33 16.29 $\pm$ 4.35 42.16 $\pm$ 7.56
NICE 92.17 $\pm$ 2.78 82.10 $\pm$ 5.55 65.65 $\pm$ 8.03 19.40 $\pm$ 5.83 46.26 $\pm$ 10.42
MLP 87.67 $\pm$ 4.41 72.58 $\pm$ 7.62 53.52 $\pm$ 9.45 11.81 $\pm$ 4.33 32.74 $\pm$ 9.24
ATM-S (Ours) 93.89 $\pm$ 1.97 85.38 $\pm$ 3.95 71.63 $\pm$ 6.17 22.84 $\pm$ 5.87 52.22 $\pm$ 8.34
@@
EEGITNet 77.14 $\pm$ 4.14 55.29 $\pm$ 6.22 31.88 $\pm$ 6.20 2.93 $\pm$ 1.70 13.00 $\pm$ 4.45
EEGConformer 79.49 $\pm$ 3.91 59.31 $\pm$ 5.83 36.67 $\pm$ 5.35 4.11 $\pm$ 1.83 17.04 $\pm$ 4.21
ShallowFBCSPNet 75.14 $\pm$ 4.45 53.06 $\pm$ 6.70 31.23 $\pm$ 6.56 3.19 $\pm$ 2.17 13.01 $\pm$ 5.15
EEGNetV4 82.60 $\pm$ 3.17 64.28 $\pm$ 5.44 42.24 $\pm$ 6.10 6.13 $\pm$ 2.40 21.23 $\pm$ 5.19
B.D. 81.49 $\pm$ 3.52 62.35 $\pm$ 6.44 40.53 $\pm$ 7.19 6.16 $\pm$ 2.40 20.45 $\pm$ 5.40
NICE 81.85 $\pm$ 2.53 63.57 $\pm$ 4.46 41.86 $\pm$ 4.68 6.43 $\pm$ 1.46 21.39 $\pm$ 3.62
MLP 80.49 $\pm$ 2.61 62.09 $\pm$ 3.88 40.37 $\pm$ 3.71 5.67 $\pm$ 1.09 19.97 $\pm$ 3.02
ATM-S (Ours) 82.88 $\pm$ 5.18 65.03 $\pm$ 8.51 44.86 $\pm$ 9.85 8.04 $\pm$ 3.34 24.62 $\pm$ 8.17
ATM-E (Ours) 83.30 $\pm$ 3.92 65.80 $\pm$ 6.98 44.84 $\pm$ 7.76 7.47 $\pm$ 2.84 23.75 $\pm$ 6.86
@@
EEGITNet 76.69 $\pm$ 12.97 56.98 $\pm$ 16.31 36.35 $\pm$ 15.11 5.75 $\pm$ 3.62 18.14 $\pm$ 9.40
EEGConformer 76.17 $\pm$ 13.13 56.29 $\pm$ 16.70 34.72 $\pm$ 14.79 3.98 $\pm$ 2.80 17.10 $\pm$ 9.21
ShallowFBCSPNet 74.32 $\pm$ 12.14 53.97 $\pm$ 15.81 33.48 $\pm$ 14.35 6.10 $\pm$ 4.61 16.53 $\pm$ 9.94
EEGNetV4 92.81 $\pm$ 2.22 83.15 $\pm$ 4.20 67.81 $\pm$ 6.11 19.51 $\pm$ 5.19 48.99 $\pm$ 6.75
B.D. 78.42 $\pm$ 8.81 58.24 $\pm$ 12.13 37.97 $\pm$ 11.38 5.88 $\pm$ 3.49 18.61 $\pm$ 7.81
NICE 92.73 $\pm$ 2.75 83.26 $\pm$ 5.47 67.96 $\pm$ 8.31 19.32 $\pm$ 5.33 49.26 $\pm$ 9.69
MLP 83.09 $\pm$ 2.54 66.70 $\pm$ 4.16 45.43 $\pm$ 4.58 7.23 $\pm$ 1.66 25.14 $\pm$ 3.66
@@
EEGITNet 77.47 $\pm$ 3.75 55.75 $\pm$ 6.01 33.60 $\pm$ 6.16 3.80 $\pm$ 2.01 14.02 $\pm$ 3.89
EEGConformer 68.02 $\pm$ 7.75 44.24 $\pm$ 9.25 23.84 $\pm$ 7.65 1.66 $\pm$ 1.37 8.62 $\pm$ 4.20
ShallowFBCSPNet 76.40 $\pm$ 4.64 54.81 $\pm$ 7.09 32.37 $\pm$ 7.28 2.49 $\pm$ 1.57 13.24 $\pm$ 5.98
EEGNetV4 82.60 $\pm$ 3.17 64.28 $\pm$ 5.44 42.24 $\pm$ 6.10 6.13 $\pm$ 2.40 21.23 $\pm$ 5.19
B.D. 85.30 $\pm$ 6.24 69.87 $\pm$ 10.09 50.82 $\pm$ 11.74 10.57 $\pm$ 5.26 29.74 $\pm$ 11.28
NICE 83.75 $\pm$ 3.21 65.82 $\pm$ 5.93 44.41 $\pm$ 6.12 7.04 $\pm$ 2.83 23.55 $\pm$ 4.86
MLP 80.49 $\pm$ 2.61 62.09 $\pm$ 3.88 40.37 $\pm$ 3.71 5.67 $\pm$ 1.09 19.97 $\pm$ 3.02
ATM-S (Ours) 87.36 $\pm$ 3.97 72.80 $\pm$ 7.02 53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
ATM-E (Ours) 87.41 $\pm$ 3.07 72.85 $\pm$ 5.51 53.15 $\pm$ 6.52 11.12 $\pm$ 3.26 32.61 $\pm$ 6.78

### v5

@@ 0.G.1 Accuracy for time windows
EEGITNet
83.23 $\pm$ 2.77
65.67 $\pm$ 3.84
43.87 $\pm$ 3.78
7.79 $\pm$ 1.27
22.12 $\pm$ 2.88
EEGConformer
89.18 $\pm$ 2.88
75.49 $\pm$ 5.30
55.32 $\pm$ 6.88
11.29 $\pm$ 3.55
33.49 $\pm$ 6.78
ShallowFBCSPNet
84.29 $\pm$ 2.56
67.13 $\pm$ 3.61
46.07 $\pm$ 4.55
8.26 $\pm$ 2.95
25.87 $\pm$ 5.03
EEGNetV4
89.03 $\pm$ 2.86
75.40 $\pm$ 5.18
56.77 $\pm$ 6.60
13.29 $\pm$ 3.99
35.50 $\pm$ 7.14
B.D.
91.14 $\pm$ 2.45
79.59 $\pm$ 4.59
62.62 $\pm$ 6.33
16.29 $\pm$ 4.35
42.16 $\pm$ 7.56
NICE
92.17 $\pm$ 2.78
82.10 $\pm$ 5.55
65.65 $\pm$ 8.03
19.40 $\pm$ 5.83
46.26 $\pm$ 10.42
MLP
87.67 $\pm$ 4.41
72.58 $\pm$ 7.62
53.52 $\pm$ 9.45
11.81 $\pm$ 4.33
32.74 $\pm$ 9.24
ATM-S (Ours)
93.89 $\pm$ 1.97
85.38 $\pm$ 3.95
71.63 $\pm$ 6.17
22.84 $\pm$ 5.87
52.22 $\pm$ 8.34
@@
EEGITNet
77.14 $\pm$ 4.14
55.29 $\pm$ 6.22
31.88 $\pm$ 6.20
2.93 $\pm$ 1.70
13.00 $\pm$ 4.45
EEGConformer
79.49 $\pm$ 3.91
59.31 $\pm$ 5.83
36.67 $\pm$ 5.35
4.11 $\pm$ 1.83
17.04 $\pm$ 4.21
ShallowFBCSPNet
75.14 $\pm$ 4.45
53.06 $\pm$ 6.70
31.23 $\pm$ 6.56
3.19 $\pm$ 2.17
13.01 $\pm$ 5.15
EEGNetV4
82.60 $\pm$ 3.17
64.28 $\pm$ 5.44
42.24 $\pm$ 6.10
6.13 $\pm$ 2.40
21.23 $\pm$ 5.19
B.D.
81.49 $\pm$ 3.52
62.35 $\pm$ 6.44
40.53 $\pm$ 7.19
6.16 $\pm$ 2.40
20.45 $\pm$ 5.40
NICE
81.85 $\pm$ 2.53
63.57 $\pm$ 4.46
41.86 $\pm$ 4.68
6.43 $\pm$ 1.46
21.39 $\pm$ 3.62
MLP
80.49 $\pm$ 2.61
62.09 $\pm$ 3.88
40.37 $\pm$ 3.71
5.67 $\pm$ 1.09
19.97 $\pm$ 3.02
ATM-S (Ours)
82.88 $\pm$ 5.18
65.03 $\pm$ 8.51
44.86 $\pm$ 9.85 8.04 $\pm$ 3.34 24.62 $\pm$ 8.17
ATM-E (Ours) 83.30 $\pm$ 3.92 65.80 $\pm$ 6.98
44.84 $\pm$ 7.76
7.47 $\pm$ 2.84
23.75 $\pm$ 6.86
@@
EEGITNet
76.69 $\pm$ 12.97
56.98 $\pm$ 16.31
36.35 $\pm$ 15.11
5.75 $\pm$ 3.62
18.14 $\pm$ 9.40
EEGConformer
76.17 $\pm$ 13.13
56.29 $\pm$ 16.70
34.72 $\pm$ 14.79
3.98 $\pm$ 2.80
17.10 $\pm$ 9.21
ShallowFBCSPNet
74.32 $\pm$ 12.14
53.97 $\pm$ 15.81
33.48 $\pm$ 14.35
6.10 $\pm$ 4.61
16.53 $\pm$ 9.94
EEGNetV4
92.81 $\pm$ 2.22
83.15 $\pm$ 4.20
67.81 $\pm$ 6.11
19.51 $\pm$ 5.19
48.99 $\pm$ 6.75
B.D.
78.42 $\pm$ 8.81
58.24 $\pm$ 12.13
37.97 $\pm$ 11.38
5.88 $\pm$ 3.49
18.61 $\pm$ 7.81
NICE
92.73 $\pm$ 2.75
83.26 $\pm$ 5.47
67.96 $\pm$ 8.31
19.32 $\pm$ 5.33
49.26 $\pm$ 9.69
MLP
83.09 $\pm$ 2.54
66.70 $\pm$ 4.16
45.43 $\pm$ 4.58
7.23 $\pm$ 1.66
25.14 $\pm$ 3.66
@@
EEGITNet
77.47 $\pm$ 3.75
55.75 $\pm$ 6.01
33.60 $\pm$ 6.16
3.80 $\pm$ 2.01
14.02 $\pm$ 3.89
EEGConformer
68.02 $\pm$ 7.75
44.24 $\pm$ 9.25
23.84 $\pm$ 7.65
1.66 $\pm$ 1.37
8.62 $\pm$ 4.20
ShallowFBCSPNet
76.40 $\pm$ 4.64
54.81 $\pm$ 7.09
32.37 $\pm$ 7.28
2.49 $\pm$ 1.57
13.24 $\pm$ 5.98
EEGNetV4
82.60 $\pm$ 3.17
64.28 $\pm$ 5.44
42.24 $\pm$ 6.10
6.13 $\pm$ 2.40
21.23 $\pm$ 5.19
B.D.
85.30 $\pm$ 6.24
69.87 $\pm$ 10.09
50.82 $\pm$ 11.74
10.57 $\pm$ 5.26
29.74 $\pm$ 11.28
NICE
83.75 $\pm$ 3.21
65.82 $\pm$ 5.93
44.41 $\pm$ 6.12
7.04 $\pm$ 2.83
23.55 $\pm$ 4.86
MLP
80.49 $\pm$ 2.61
62.09 $\pm$ 3.88
40.37 $\pm$ 3.71
5.67 $\pm$ 1.09
19.97 $\pm$ 3.02
ATM-S (Ours)
87.36 $\pm$ 3.97
72.80 $\pm$ 7.02
53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
ATM-E (Ours) 87.41 $\pm$ 3.07 72.85 $\pm$ 5.51
53.15 $\pm$ 6.52
11.12 $\pm$ 3.26
32.61 $\pm$ 6.78